OpenStudy (anonymous):

4 over 27= x over 18 what is the value of x

mathslover (mathslover):

\[\large{\frac{4}{27}=\frac{x}{18}}\] \[\large{\frac{4}{27}\times 18=x}\] \[\large{\frac{4}{\cancel{27}^{3}}*\cancel{18}^2=x}\] \[\large{\frac{4}{3}*2=x}\] \[\large{\frac{8}{3}=x}\]

mathslover (mathslover):

@EmoLover3 got it?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yep

mathslover (mathslover):

nice : sorry for direct solution didn't have time and thought u needed soln hence posted it please prefer to learn it step by step best of luck!!!

OpenStudy (anonymous):

:D